OpenStudy (anonymous):

How do I solve: 4 = (x + 1/2)^2?

OpenStudy (anonymous):

well.....take a square root on both sides. x + 1/2 = +/- 2. So, x = 3/2 or -5/2

OpenStudy (anonymous):

so... 4 will be +/- 2, and then x + 1/2?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

x + 1/2 = 2 OR x + 1/2 = -2 x = 3/2 OR x = -5/2

OpenStudy (anonymous):

but why is it (x + 1/2)^2?

OpenStudy (anonymous):

Is it not what is given in the problem?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ohhh, i get it. Thank you!